OpenStudy (anonymous):

Which of the following correctly completes the square for the equation below? x^2+6x=5 A) (x+3)^2=5 B) (x+3)^2=14 C) (x+6)^2=5 D) (x+6)^2=14 Please help me...

OpenStudy (helder_edwin):

to complete the square you have to do this \[ \large x^2+6x+a=5+a \] now since \[ \large (y+z)^2=y^2+2yz+z^2 \] what is the value of a?

OpenStudy (anonymous):

Is it one?

OpenStudy (helder_edwin):

no! watch this \[ \large (x+k)^2=x^2+2xk+k^2 \] compare with what you have

OpenStudy (anonymous):

To be honest I don't know what I'm doing.. like at all.

OpenStudy (helder_edwin):

if you compare the last i wrote with what u have then \[ \large 6x=2kx \] then 6=2k then k=3 OK with this a=?